OpenStudy (anonymous):

-7/10=k-6.1/10 i need lots of help plz

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

@Abhisar

OpenStudy (sleepyjess):

Are you trying to solve for k?

OpenStudy (anonymous):

yes

OpenStudy (sleepyjess):

Ok, what do you think you will do first?

OpenStudy (anonymous):

i forgot and i cant go look back

OpenStudy (sleepyjess):

Ok, personally, I like to change fractions into a decimal.

OpenStudy (anonymous):

ok

OpenStudy (sleepyjess):

Is it \(-\dfrac{7}{10}=\dfrac{k-6.1}{10}\) or \(-\dfrac{7}{10}=k-\dfrac{6.1}{10}\)

OpenStudy (anonymous):

no i will copy it if i can

OpenStudy (sleepyjess):

Could you do a screenshot?

OpenStudy (anonymous):

−  7 ___ 10 = k − 6 1 ___ 1

OpenStudy (sleepyjess):

So it is like the first one I put up?